OpenStudy (anonymous):

Rewrite the equation in Ax+By = C form. Use integers for A, B, and C y+6 = 3(x-2)

OpenStudy (anonymous):

first release the brackets: y+6=3x-6 now subtract 6 from both sides y+6-6=3x-6-6 y=3x-12

OpenStudy (anonymous):

now subtract 3x from both sides y-3x=-12+3x-3x y-3x=-12